Operations Report – January 13th

  

 Operations Report

Crew 219 Operations Report 13-JAN-2020

Sol: 8

Name of person filing report: Alejandro Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 1.5

From what time last night: 18:00

To what time this morning: *failure stamp 19:30

List of any additional daytime hours when the generator was run:

Start: 06:00

Tuned off: 15:00

Hours run: 9

Solar- SOC% (Before the generator is run at night): 83%

06:30 = 10%; 07:00 = 14%; 08:30 = 22%; 10:00 = 33%; 11:45 = 46%; 12:30
= 51%; 13:30 = 55%; 14:30 = 100%; 15:30 = 98%; 16:30 = 90%; 17:30 =
85%; 18:00 = 83%; 18:30 = 81%;

Diesel Reading - 60%

Station Propane Reading- 39%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0147647,2

Water (static tank) (gal) – 384

Static to loft Pump used –Yes

Water in Green Hab (gal): 166.21

Water in Science Dome (gal): 0

Toilet Tank Emptied –Yes

Diemos rover used: Still in the Work shop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- Yes

Starting Hours: 124.9

Beginning Charge: 100%

Ending Hours: 125.4

Ending Charge: 27%

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

Starting Hours: 130.8

Beginning Charge: 100%

Ending Hours: 131.3

Ending Charge: 38%

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where?

CrewCar used and why, where?

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: Nominal

Summary of Hab: generator failed due to underspeed issue

Summary of Science Dome operations: Temperatures are hovering around 4
°C during the day and 0 °C during the night.

Summary of Ram operations: Power converter began a high-pitched
whining noise during the day. Outpost and CO assessed the problem.
Outpost advised to include it in the notes of the operation report.

Summary of any observatory issues: Nominal

Summary of Health and safety issues: See HSO report

Questions, concerns, and requests to Mission Support: None. 

Operations Report – January 13th

  

 Operations Report

Crew 219 Operations Report 13-JAN-2020

Sol: 8

Name of person filing report: Alejandro Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 1.5

From what time last night: 18:00

To what time this morning: *failure stamp 19:30

List of any additional daytime hours when the generator was run:

Start: 06:00

Tuned off: 15:00

Hours run: 9

Solar- SOC% (Before the generator is run at night): 83%

06:30 = 10%; 07:00 = 14%; 08:30 = 22%; 10:00 = 33%; 11:45 = 46%; 12:30
= 51%; 13:30 = 55%; 14:30 = 100%; 15:30 = 98%; 16:30 = 90%; 17:30 =
85%; 18:00 = 83%; 18:30 = 81%;

Diesel Reading - 60%

Station Propane Reading- 39%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0147647,2

Water (static tank) (gal) – 384

Static to loft Pump used –Yes

Water in Green Hab (gal): 166.21

Water in Science Dome (gal): 0

Toilet Tank Emptied –Yes

Diemos rover used: Still in the Work shop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- Yes

Starting Hours: 124.9

Beginning Charge: 100%

Ending Hours: 125.4

Ending Charge: 27%

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

Starting Hours: 130.8

Beginning Charge: 100%

Ending Hours: 131.3

Ending Charge: 38%

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where?

CrewCar used and why, where?

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: Nominal

Summary of Hab: Generator had a power malfunction causing no charge
with the engine running. SOC became under leveled before 06:00 and
safety mode was initiated.

Summary of Science Dome operations: Temperatures are hovering around 4
°C during the day and 0 °C during the night.

Summary of Ram operations: Power converter began a high-pitched
whining noise during the day. Outpost and CO assessed the problem.
Outpost advised to include it in the notes of the operation report.

Summary of any observatory issues: Nominal

Summary of Health and safety issues: See HSO report

Questions, concerns, and requests to Mission Support: None. 

Sol Summary – January 13th

  

 Crew 219 Sol Summary Report 13-01-2020

Sol: 8

Summary Title: Oh, Mondays...

Author's name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: The crew was woken abruptly around 0600 to a
generator failure and low power, and immediately responded according
to guidance from Outpost. We then boiled water on the stove for
coffee, and got started with the days work. Some worked on science
while others transplanted plants in the GreenHab. H. Blackburn made a
lunch of chili and rice before EVA #13 pushed out towards Beige Moon,
and those left behind continued their tasks. EVA #13 collected their
samples, but suffered an H&S-related incident (see HSO report) forcing
their return by 1508. K. Crisman resolved the situation swiftly,
decisively, and admirably. With the whole crew back at nominal status,
A. Elnajdi and R. Raphael prepared an incredible dinner for the crew
while everyone finished and proofread reports for this evening’s Comms
window.

Look Ahead Plan: One EVA, continued sample analysis, and maintenance
as necessary

Anomalies in work: Generator Failure

Weather: Early morning fog, cold and sunny all day

Crew Physical Status: Nominal at present (See HSO Report)

EVA:

EVA #13 – Geologic Sample Collection around Beige Moon (0516000/4254300)

Reports to be filed:

Sol Summary

EVA Report #13

EVA Request #14

HSO Report

Operations Report

GreenHab Report

Research Report

Journalist's Report

Photos (Including Photo of the Day)

Support Requested:

-Quart-sized ziploc bags

-Medical tape

-Refer to Operations Report for maintenance-related request 

EVA Report – January 13th

  

 Crew 219 EVA Report 13 JAN 2020

EVA#: 13

Author: Nathan Hadland

Purpose of EVA: The primary purpose of this EVA was sampling for our
geology project. Post-EVA, we collected data for our biometrics
project.

Start time: 1313 hrs

End time: 1513 hrs

Narrative: At 1313 hours, EVA #13 initiated a 5-minute decompression
and moved to Curiosity and Spirit. After reporting SOC and hours, we
progressed north along Cow Dung Road until we reached Brahe Hwy. We
moved west on foot through White Moon. After taking 5 regolith
samples, the crew then continued west to Beige Moon. The crew then
took 5 additional samples and then turned around and headed east back
to the rovers. Curiosity and Spirit arrived back at the Hab at 1508
hours. There was a health and safety incident at the end of the EVA;
refer to the HSO report for more details.

Destination: Beige Moon

Coordinates: 0516000/4254300

Participants: Nathan Hadland (XO), Alejandro Perez (ENG), Robinson
Raphael (ASTRO), Abdul Elnagdi (GEO)

Roads and routes per MDRS Map: North on Cow Dung Road (0110) parking
at the intersection of Brahe Hwy (1575) and walking west through White
Moon and Beige Moon along Brahe Hwy.

Research Report – January 13th

  

 Crew 219 Science Report 13-JAN-20
Crew Science Officer: Hannah Blackburn

1. Biometrics and Neurobehavioral Research

We continued taking sleep logs and post-EVA biometrics and surveys.

2. UAV

Nothing to report.

3. Dust Mitigation for Optical Mirrors

Cleaned the optical mount using air pump.

4. Astrophotography of Celestial Bodies

False color image of the Crab Nebula will be attached with the next
Astronomy Report.

5. Remediation of Mars Regolith

Samples observed under microscope.

6. Chemical and Mineralogical Composition of the MDRS Site

Samples were dried in the oven using glass beakers. Additional samples
collected during EVA 13.

7. Protocols for the Discovery of Life on Mars

Nothing to report.

Glassware check out:

10 glass beakers (sizes between 200 and 500 ml) were used to dry
regolith samples.

A reminder to all crewmembers: There's a $300 fine for using any
glassware material without MDRS Mission support permission.

Journalist Report – January 13th

  

Crew 219 JOURNALIST REPORT 13JAN2020

AUTHOR: Cynthia Montanez

SOL 8

AVANTGARDE

As darkness began to fade this morning, a distant sound awakens Crew
219. Confused as to what the sound was, the HSO (Keith Crisman) found
the nearest radio and asked “Say again?” With a reply from Outpost
stating the low SOC percentage, Crew 219 finds themselves in safety
mode. While in safety mode, each crew member unplugged all their
devices and appliances within the station to conserve the remaining
energy.

At 10:00 A.M., LSO (Hannah Blackburn) and I decided to make the best
out of this situation by having a Martian Spa Day. After an hour of
detangling my curly hair, the LSO and I set up a mini station by the
airlock for their spa. The spa consisted of a detangling brush, Head
and Shoulder shampoo and conditioner, towel, and a dish bin filled
with warm water. It was an experience we both enjoyed. Not to mention,
the results that came from our Martian Spa was beautiful clean hair.

By about 11 am, Crew 219 was able to turn on their lights. With
everyone celebrating, I decided to head to the GreenHab for my first
temperature reading of the day and a quick watering. While watering,
the CO (Dave Masaitis) entered and assisted by transplanting some aloe
vera succulents and cucumbers. We spent time in the GreenHab
conversing about planting methods, future plans, etc. It was a brief
discussion, but it was one that brought us closer together.

As the rest of the day played out, while part of the crew was out on
EVA, the rest of crew 219 found ways to keep themselves busy by either
cooking, reading a good book, working in the lab, or conversing with
one another.

Overall, today was a good day on white Mars. 

GreenHab Report – January 13th

  

 Crew 219 GreenHab Report 13-JAN-20

Crew GreenHab Officer: Cynthia Montanez

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 25.2 °C; 17.7 %

Average Max: 29.3 °C, 20.7%

Average Min: 19.6 °C, 15%

11:30 A.M.

Floor Unit: 21 °C

Electronic: 27.8 °C

humidity 17%

Max: 27.1 °C; 24%

Min: 13.9 °C; 16%

3:00 P.M.

Floor Unit: 15 °C

Electronic: 26.2 °C

humidity 17%

Max: 31.6 °C; 18%

Min: 24.7 °C; 14%

5:00 P.M.

Floor Unit: 16 °C

Electronic: 21.8 °C

humidity 19%

Max: 29.2 °C; 20%

Min: 20.3 °C; 15%

Hours of supplemental light: Light system 7:00PM-12:00AM.

Daily water usage for crops: 8.65 Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – 166.21 Gal.

Time(s) of watering for crops:

11:30 A.M.

SEASONAL PLANTING HISTORY:

Change to crops: N/A

OTHER NOTES:

1st sprouts: N/A

Harvest: At 11:50 A.M., 4 grams of sage was harvested.

Other: Two aloe vera plants and two cucumber plants were transplanted
between 11:35 A.M. – 11:50 A.M.